three years ago the fairchild co issued 20 year, 7.75 percent semiannual coupon bonds at par. today the bonds are quoted at 102.6. what is the firm's pretax cost of debt
Answer #1

Suppose the face value (or par value) of the bond be $1000
Given that the bonds are quoted at 102.6, it means present value is 102.6%*1000=1026
Coupon rate is 7.75%
As the coupon payments are made semiannually, the semi annual coupon rate=7.75%/2=0.03875

Semiannual coupon payments=(Semiannual coupon rate)*(Face value)
=(0.03875)*(1000)=38.75
Remaining time period =20-3=17 years
Number of periods when the interest is paid semiannually is 17*2=34 years

Pre tax cost of debt refers to yield to maturity or YTM, we can determine the ytm using excel.


Now, the semiannual yield to maturity=3.74%
Annualized yield to maturity=3.74%*2=7.48%

Answer: Pre tax cost of debt is 7.48%